OpenStudy (xapproachesinfinity):

how to prove 5657 is prime

OpenStudy (empty):

Divide it by all natural numbers less than \(\sqrt{5657}\) and see that there are no whole number solutions. Easy if you have a computer I guess, but difficult by hand.

OpenStudy (xapproachesinfinity):

i need mathematical proof, i was given this on exam i just skipped it i was trying this at the beginning 5656<5657<5658 5656 and 5658 are both composite but didn't know how to carry the proof

OpenStudy (empty):

There's nothing invalid about what I said, it constitutes mathematical proof and takes less than a second with a computer. But I understand that's not what you want. The problem is, what kind of proof are you looking for?

OpenStudy (xapproachesinfinity):

question asks for any mathematical proof

OpenStudy (wolf1728):

Empty, actually to save a little time you have to divide the prime being tested by all PRIME numbers (2,3,5,7,11, 13, etc).

OpenStudy (xapproachesinfinity):

why do you want to use numbers less that sqrt(5657)

OpenStudy (xapproachesinfinity):

that's way too much of work to test on all smaller primes

OpenStudy (xapproachesinfinity):

this is a hand problem not a machine problem

OpenStudy (empty):

True, I guess if I were to do it, I'd use a sieve and it would automatically skip out on those, good catch @wolf1728 @xapproachesinfinity The reason you do that is cause the smallest number of numbers a number can be divisible by and not be prime is two, so basically we're testing to see if it can be possibly written as the product of two numbers... (sorry that's all super wordy sounding, I'll just write this) 5657=a*b So either a=b or a<b, there's no reason to test them all. I'll give an example for trying to see if 9 is prime: Test 9/2 = not an integer Test 9/3 = an integer (let's keep going though) Test 9/4 = not an integer but notice, that 9/4 < 3 and we already tested all the numbers up to 3, so we're wasting our time by checking numbers larger than the square root of 9... Hopefully this kinda makes sense.

OpenStudy (empty):

I have no idea how to show this is prime, but I know 'how' to show that it's prime.

OpenStudy (xapproachesinfinity):

i got that !

OpenStudy (xapproachesinfinity):

so how would we save time and get to the right numbers closest to sqrt(5657)

OpenStudy (empty):

This is impossible on a test, your teacher is looking for something else no doubt in my mind

OpenStudy (empty):

Does anything on here look like something you've seen before? Lucas primality test? (never heard of this before so I don't know) https://en.wikipedia.org/wiki/Primality_test#Number-theoretic_methods

OpenStudy (xapproachesinfinity):

i have no idea, he stated show by any mathematical method

OpenStudy (xapproachesinfinity):

no ihaven't studied number theory

OpenStudy (empty):

What class is this? Are you supposed to prove that it's prime or just show an algorithm that could prove that it's prime?

OpenStudy (dan815):

ya id do the same thing check all primes under sqrt5657

OpenStudy (dan815):

primes under 75 thats not too bad

OpenStudy (dan815):

ill think if there is some neat way to show prime without checking hmm

OpenStudy (dan815):

u know the inverted circle stuff?

OpenStudy (xapproachesinfinity):

hmm the prof left he didn't explain everything hahha

OpenStudy (dan815):

if we can show this number is part of some sequence where only primes are being generated

OpenStudy (dan815):

okay well this inverted circle thing i found on youtube about how the radiuses of these circles in a particular geometry all had prime denominators

OpenStudy (xapproachesinfinity):

hmm i see

OpenStudy (dan815):

something like that maybe.. lets see if we can think of some seuqnece that we can prove must only contain primes and that this prime is in it

OpenStudy (dan815):

we can prove if a number is not prime from fermats little theorem how aobut that

OpenStudy (dan815):

Calculate 2^(n-1) mod n. If you don't get 1 mod n, then n is not prime.

OpenStudy (xapproachesinfinity):

oh yeah sound much better care to show it please

OpenStudy (dan815):

okay so after finding what 2^(n-1) is you would do the chinese remainder theorem to see the GCD

OpenStudy (dan815):

and ull end up with GCD = 1

OpenStudy (dan815):

and why its working, ud have to understand fermats little theorem

OpenStudy (xapproachesinfinity):

ok

OpenStudy (dan815):

https://www.youtube.com/watch?v=w0ZQvZLx2KA

OpenStudy (xapproachesinfinity):

seems a straight forward theorem

OpenStudy (dan815):